# Assessing the Uncertainty and Robustness of Object Detection Models for Detecting Stickers on Laptops

### v1

@@ 8 Related Works
We discuss related works from two perspectives, i.e., UQ in DNNs and robustness assessment of DNNs.
UQ in DNNs. UQ helps enhance model reliability and trustworthiness. Various UQ methods have been studied to measure confidence in the model predictions. BNN provides a probabilistic framework for DNNs through Bayesian inference [5]. It is systematic but computationally expensive and has a complex implementation. As a Bayesian approximation, MC-Dropout is a practical and widely used UQ method [6], which reduces the computational cost. Deep Ensembles (DE) [44] involves independently training multiple DNNs with different initialization, and the variance in the predictions of the multiple DNNs serves as an uncertainty measure. DE is robust and reliable in uncertainty estimation but computationally and memory-wise expensive, requiring training and storing multiple models. UQ methods have been applied in many domains, e.g., cyber-physical systems [45, 46], computer vision tasks [47, 19, 48], and heathcare [49]. In contrast, this paper adopts MC-dropout as the UQ method to quantify the uncertainty of sticker detection tasks in laptop refurbishing robots, thereby studying their real-world application in a new context.
Robustness Assessment of DNNs. Robustness measures DNNs’ reliability. Recent robustness evaluation methods focus on adversarial robustness [7], which refers to the ability of DNNs to maintain performance and provide reliable outputs in the face of various perturbations, noises, and adversarial attacks. For instance, Carlini and Wagner [4] constructed three adversarial attack methods to evaluate the robustness of defensively distilled networks. Madry et al. [8] proposed a Projected Gradient Descent (PGD) attack to assess the adversarial robustness of DNNs from the robust optimization perspective. PGD employs testing accuracy as an estimation to evaluate the model’s robustness. Croce and Hein [50] further proposed two extensions of the PGD attack to address the problem of overestimating robustness in PGD. Yu et al. [51] qualitatively interpreted the adversarial vulnerability and robustness of DNNs and proposed a quantitative metric for the model-intrinsic robustness evaluation by measuring the model’s prediction divergence. Brendel et al. [52] combined decision-based boundary attack with the gradient-based adversarial attack to achieve fast and accurate robustness evaluation. This paper employs a dense adversarial attack technique to measure adversarial robustness. Regarding the robustness metrics, in addition to measuring the robustness using the performance metric (i.e., mAP), we further calculate the robustness score concerning UQMs.
@@ 9 Conclusions and Future Work
In this paper, we conduct an empirical study to evaluate the detection accuracy, prediction uncertainty, and adversarial robustness of six sticker detection models in the laptop refurbishment process. We adopt the Monte Carlo Dropout method to quantify the prediction uncertainty and measure uncertainty from two aspects: uncertainty in classification and regression. In addition, we present novel robustness metrics to evaluate the robustness of the models regarding detection accuracy and uncertainty. As for the benchmark datasets, in addition to the original dataset, we synthesize two datasets using two vision language models. Besides, we generate three adversarial datasets using an adversarial attack method. The results show different performances of different models, and Faster R-CNN_v2 and RetinaNet_v2 achieve the overall best performance in terms of all metrics. We also provide model selection guidelines and lessons learned. Future works include quantifying uncertainties of the overall laptop refurbishing process and investigating methods to generate large-scale realistic laptop image datasets.
Acknowledgments
The work presented here is supported by the RoboSapiens project funded by the European Commission’s Horizon Europe programme under grant agreement number 101133807.

### v2

@@ 8 Related Works
UQ in DNNs. UQ helps enhance model reliability and trustworthiness. Various UQ methods have been studied to measure confidence in the model predictions. BNN provides a probabilistic framework for DNNs through Bayesian inference [7]. It is systematic but computationally expensive and has a complex implementation. As a Bayesian approximation, MC-Dropout is a practical and widely used UQ method [3], which reduces the computational cost. Deep Ensembles (DE) [39] involves independently training multiple DNNs with different initialization, and the variance in the predictions of the multiple DNNs serves as an uncertainty measure. DE is robust and reliable in uncertainty estimation but computationally and memory-wise expensive, requiring training and storing multiple models. UQ methods have been applied in many domains, e.g., cyber-physical systems [40, 41], computer vision tasks [42, 15, 43], and healthcare [44]. In contrast, this paper adopts MC-dropout as the UQ method to quantify the uncertainty of sticker detection tasks in laptop refurbishing robots, thereby studying their real-world application in a new context.
Robustness Assessment of DNNs. Robustness measures DNNs’ reliability. Recent robustness evaluation methods focus on adversarial robustness [45, 46, 47], which refers to the ability of DNNs to maintain performance and provide reliable outputs in the face of various perturbations, noises, and adversarial attacks. For instance, Carlini and Wagner [2] constructed three adversarial attack methods to evaluate the robustness of defensively distilled networks. Madry et al. [19] proposed a Projected Gradient Descent (PGD) attack to assess the adversarial robustness of DNNs from the robust optimization perspective. This paper employs a dense adversarial attack technique to measure adversarial robustness. Regarding the robustness metrics, in addition to measuring the robustness using the performance metric (i.e., mAP), we further calculate the robustness score concerning UQMs.
@@ 9 Conclusions and Future Work
In this paper, we conduct an empirical study to evaluate the detection accuracy, prediction uncertainty, and adversarial robustness of six sticker detection models in the laptop refurbishing software. We adopt the Monte Carlo Dropout method to quantify the prediction uncertainty and measure uncertainty from two aspects: uncertainty in classification and regression. Besides, we present novel robustness metrics to evaluate the robustness of the models regarding detection accuracy and uncertainty. The results show Faster R-CNN_v2 and RetinaNet_v2 achieve the overall best performance regarding all metrics. Future works include quantifying uncertainties of the entire laptop refurbishing process and automatically generating realistic image datasets.
Acknowledgement
The work is supported by the RoboSapiens project funded by the European Commission’s Horizon Europe programme under grant agreement number 101133807. Jiahui Wu is also partially supported by the Co-tester project from the Research Council of Norway (No. 314544).